OpenStudy (anonymous):

In analyzing the income statement of Bob Compay, cost of goods sold decreased from 2013 to 2014 by 8.2%. The cost of goods sold was $19,000 in 2014. The cost of goods sold in the nearsest cent in 2013 was:

OpenStudy (anonymous):

let the cost of good be x

OpenStudy (anonymous):

so in 2014 it will be \[x - \left(\begin{matrix}8.2 \\ 10?\end{matrix}\right)x\]

OpenStudy (anonymous):

so this equation will be equal to 19000

OpenStudy (anonymous):

can u find it now

OpenStudy (anonymous):

there ??

OpenStudy (anonymous):

Not yet

OpenStudy (anonymous):

did u understand how did i take x

OpenStudy (anonymous):

Yes..I came up with $18,999.18 but not sure this is correct

OpenStudy (anonymous):

noww u do it it will be correfct

OpenStudy (anonymous):

and \[x -\frac{ 82 }{ 1000 }x\]

OpenStudy (anonymous):

this will be equal to 19000

OpenStudy (anonymous):

Ok I see that

OpenStudy (anonymous):

yeah u see it and tel me

OpenStudy (anonymous):

$20,679.71

OpenStudy (anonymous):

no its 1742 i calculated it

OpenStudy (anonymous):

17442*

OpenStudy (anonymous):

can i get a medal :O